### v1

# The Hot Jupiter Radius Anomaly and Stellar Connections

Abstract
The extremely close proximity of hot Jupiters to their parent stars has dramatically affected both their atmospheres and interiors, inflating them to up to twice the radius of Jupiter. The physical mechanism responsible for this inflation remains unknown, though many proposals have been put forward. I will review the known hot Jupiter population, the proposed inflation mechanisms, and the evidence for and against them collected thus far. In doing so, I will cover the ways that hot Jupiter interiors may be simulated computationally in detail, and present some useful formulas for estimating their radii, heating, intrinsic temperature, and tentative magnetic field strength. I will also cover the related issues of hot Jupiter intrinsic temperatures and radiative-convective boundaries, the potential connection with planetary magnetic fields, and the effects of stellar tides on the planet. Finally, I conclude with the suggestion that more than one mechanism may be operating in concert with each other and propose various avenues for future progress in understanding these objects.
1 Introduction
The discovery of on close-in orbits around their parent stars (Mayor and Queloz 1995, the first being) came as something of a surprise and raised the question of how such objects could be formed (e.g., Pollack et al. 1996; Ida and Lin 2004). Some argued the planets could have formed in-situ (Bodenheimer et al. 2000, e.g.,), while others favored migration from wider separations due to interactions with the disk (e.g., Masset and Papaloizou 2003) or the star (eccentric migration Marzari and Weidenschilling 2002; Ford and Rasio 2006). This was the first of two great problems in physics. The second emerged when the transit method revealed a radius for HD 209458 b: (Charbonneau et al. 2000) and (Henry et al. 2000). This large size stood in stark contrast with our previous understanding of giant planet physics, which predicted a radius of (Bodenheimer et al. 2001; Guillot and Showman 2002).
Initially, this discrepancy was understood to be either the result of observational and theoretical uncertainties (Burrows et al. 2003, 2004) or as an exceptional case of a large planet (Bodenheimer et al. 2003) having an unusually hot interior due to tidal interactions with the parent star (Bodenheimer et al. 2001). Guillot and Showman (2002) proposed that while the planet indeed must be very hot (which was supported by Baraffe et al. 2003; Chabrier et al. 2004), the most likely source of heating was not tidal interactions but rather the intense instellation from the parent star heating the interior through a previously unknown physical mechanism. A small number of additional transiting hot Jupiters were discovered, including OGLE-TR-56 b (Konacki et al. 2003) and TrES-1 Alonso et al. (2004), revealing more moderately inflated hot Jupiters similar to HD 209458 b.
The number of known hot Jupiters grew substantially in the late 2000s thanks in large part to discoveries made by the Hungarian Automated Telescope (HAT; Bakos et al. 2004, 2007) and Wide Angle Search for Planets (WASP; Pollacco et al. 2006; Collier Cameron et al. 2007) surveys. These exhibited consistently larger radii than theoretical models could produce (Fig. 1), demonstrating that the radius anomaly was not a special case for the first few hot Jupiters, but instead a process broadly affecting the entire population. Subsequent discoveries using Kepler (Borucki et al. 2010), K2 (Howell et al. 2014), CoRoT (Auvergne et al. 2009), HAT-South (HATS; Bakos et al. 2013), the Transiting Exoplanet Survey Satellite (TESS; Ricker et al. 2014), and others have expanded the number of transiting hot Jupiters with mass measurements to over 500! In this chapter, I will give an overview of this population, explain how interior structure models of these objects are calculated, list some of the proposed explanations for their radii, and identify important observational implications of these hypotheses.
2 The Observed Population
The current set of giant exoplanets with known radii and masses is shown in Fig. 1, plotted as radius against flux. The black dashed line depicts the radius of a pure H/He object with no extra internal heating. As originally observed by Miller and Fortney (2011) and Demory and Seager (2011), radius inflation can be seen to turn on at erg s-1 cm-2 ( K), marked with a vertical dotted line. Above this radii are correlated with the incident flux. The most highly inflated planets reach at in the 1500-1800 K range, and are generally somewhat less massive than Jupiter. On average, more massive planets are smaller for a given incident flux.
The observed hot Jupiters population predominantly orbits F, G, and K dwarf stars due to a combination of observation bias against their detection around high-mass stars (the transit depth is much lower) and an inherently low occurrence rate around M dwarfs (Johnson et al. 2010). This variation in stellar mass and therefore luminosity allows us to examine whether the radii are most strongly correlated with the incident flux (most related to atmospheric effects) or the orbital period (more related to tidal effects). Weiss et al. (2013) finds that the incident flux is a much better predictor of the radius anomaly than the period. They also provide another useful result: an empirical mass-flux-radius relation,
for planets with . is the planetary radius in , is the planetary mass, and is the incident flux onto the planet in erg s-1 cm-1.
For completeness and comparison we’ll also mention a few other empirical fits to hot Jupiter radii. Chen and Kipping (2017) provide a rigorously-derived mass-radius relation for a wide range of planetary masses and an easy-to-use code for calculating predictions in either direction, but note that they do not use the incident flux as a predictor. They therefore output radii corresponding with a typical flux of around erg s-1 cm-1. For hot Jupiters in particular, Thorngren et al. (2021) experimented with a large number of functional forms with a larger sample size but recommends a similar form to that of Weiss et al. (2013):
where is the radius in , is the mass in , and is the incident flux in erg s-1 cm-1. Note the addition of a cross-term between flux and radius: the inflationary effect of flux decreases with increasing mass.
In some cases it may be desirable to estimate the radius in the absence of anomalous heating – e.g. to calculate the radius anomaly as in Laughlin et al. (2011). Fortney et al. (2007) calculated a grid of model radii that may be interpolated for this purpose. Alternatively, (Thorngren et al. 2019b) estimate the following for giant planets with :
where is the planetary radius in and is the planetary mass in . The subscript “uninflated” reminds readers that this is the planet radius assuming no inflation, based on the warm giant planets. Note that a predictive error is also estimated as a function of mass.
Finally, we will mention a few hot Jupiters of particular interest. HD 209458 b was the first discovered hot Jupiter, but it also is a fairly typical example of the class that is extremely well-studied, making it a good choice if readers want a single ”typical case” to study. KELT-9 b (Gaudi et al. 2017) is the hottest known hot Jupiter with an impressive K, rivaling some K stars in surface temperature and thereby hosting an atmosphere with radically different chemistry from typical hot Jupiters (e.g., Hoeijmakers et al. 2018). The largest planet known is one of WASP-17 b, WASP-78 b, or WASP-79 b, all hot Jupiters of about 2 whose radius uncertainties prevent us from knowing which is largest. WASP-12 b, WASP-121 b, and WASP-103 b are believed to be the most tidally-distorted planets (Wahl et al. 2021), though by their radius vs Roche radius WASP-19 b and Qatar-10 b may be comparably distorted. Lastly, the shortest-period hot Jupiter with well-determined properties is currently TOI-2109 b (Wong et al. 2021), orbiting its star in just 16 hours.
2.1 Sampling Biases
In this section, we have looked at the full sample of known transiting hot Jupiters, but one should bear in mind the observational biases that might be affecting it. Primarily, this refers to the biases of the transit method, whereby larger objects produce more easily detectable transit signals. Hot Jupiters are sufficiently large and short-period that for bright stars they are extremely likely to be successfully observed, limiting the importance of this bias. Indeed, Yee et al. (2021) found that for stars with , the (excluding the poles and galactic plane) is statistically nearly 100%! Further, they argue that TESS might reasonably expand this completeness to a limiting G-magnitude of about 12.5 (see also Sullivan et al. 2015; Zhou et al. 2021). This will present an exciting opportunity to better understand these planets, but we should still remember that the order in which they are discovered will not be random; the selection function for follow-up is undoubtedly complex, and unusual planets may be prioritized for confirmation.
Another characteristic that biases the observed population of hot Jupiters is the subtle dependence on the age of the parent star. The principal cause of this is (Hamer and Schlaufman 2019), in which the closest-in hot Jupiters raise tides in their host stars that saps them of their angular momentum, decaying their orbits and eventually causing (Valsecchi et al. 2014, 2015). The tidal circularization of eccentric planets into hot Jupiters is expected to be generally very fast (Jackson et al. 2008), with a few rare cases of planets being observed presently undergoing this process, such as HD 80606 b (Naef et al. 2001). Finally, while higher-mass giant planets are thought to be fairly resistant to (Murray-Clay et al. 2009; Dwivedi et al. 2019), planets near Saturn’s mass or lower can still undergo substantial XUV-driven mass loss (Caldiroli et al. 2022; Kurokawa and Nakamoto 2014; Thorngren et al. 2023). As such, hot Jupiters orbiting young stars (e.g., Newton et al. 2019; Zhou et al. 2021) are a valuable point of comparison from before these processes have completed.
3 Modeling Hot Jupiter Interiors
The radius of a giant planet is mainly determined by the thermal state and composition of its interior, with only a small portion (, depending on the temperature) of the radius contributed by the atmosphere. As such, modeling these interiors is the first step in understanding the radius inflation problem. Even very modern computers are not fast enough to run fluid-dynamical simulations in three dimensions over billion-year timescales, so the long-term evolution of giant planets is relegated to one-dimensional models. In this section, we’ll review how one may construct such models of giant planets at a single point in time (static models), how these may be evolved forward through time, and additional relevant physics that one may wish to include in modeling these objects, depending on circumstance.
It should be mentioned that there are existing software packages that can be used to model , the most prominent being MESA (Paxton et al. 2013). While originally built for modeling the interiors and evolution of stars, the fundamental PDEs being solved are the same, allowing it to be successfully adapted to gas giant use cases on numerous occasions (Chen and Rogers 2016; Mankovich et al. 2016, e.g.,). Still, prospective users should be aware that these modifications are non-trivial. Another open-source option is MAGRATHEA (Huang et al. 2022); while mainly focused on the simulation of terrestrial or sub-Neptune planets, it can also be used for the simulation of true gas giant interiors.
3.1 Static Models
At an instant in time, the state of a giant planet can be described by its total mass and a complete description of the interior composition and temperature at a given mass shell . For a one-layer fully-convective planet, the composition can be the metal mass-fraction of the planet – 0 representing a metal-free planet and 1 being a planet composed entirely of metal (no hydrogen or helium). Thanks to our assumption of a fully-convective planet, the thermal state can be described by the specific entropy . An equation of state maps the specific entropy to a temperature for any given temperature, and the curve through temperature-pressure space is known as the adiabat. For giant planets, the of a hydrogen-helium envelope ranges from 5 (very “cold”) to 12 (very “hot”) per baryon, where is the Boltzmann constant. Giant planets begin their life with specific entropy somewhere in the 8-12 per baryon range (Marley et al. 2007; Cumming et al. 2018), which then decreases as energy is released through the atmosphere.
Given the mass, composition, and temperature structure of the planet, we can calculate its radius by solving the equations of hydrostatic equilibrium, mass conservation, and an appropriate (EOS):
Eq. 6 shows the across the equations of state of multiple components with mass fraction and density as a function of pressure and specific entropy. For these, one must decide what materials are to be modeled. A typical choice would start with hydrogen and helium at the solar (see Asplund et al. 2009) or protosolar ratios. For the latter, Serenelli and Basu (2010) give and , depending on the exact assumptions made. With a middling and (Asplund et al. 2009), we have . To obtain the density of H/He one should interpolate from an EOS table containing densities at various pressures and specific entropies. Chabrier et al. (1992) was the dominant choice for over a decade, but more recent ab-initio calculations (Nettelmann et al. 2008; Militzer and Hubbard 2013; Chabrier et al. 2019) have supplanted it. Chabrier and Debras (2021) would be a strong choice for H/He as of the time of writing, as it combines information from both Militzer and Hubbard (2013) and Chabrier et al. (2019). Fig. 3 shows an example interior model for a hot and warm Jupiter, calculated using these equations and the Chabrier et al. (2019) EOS.
Gas giants interiors also contain metals, which are typically represented in the equation of state by water and/or a representative rock (often olivine, since it is a large portion of the Earth’s mantle). The outer solar system icy moons have roughly 50-50 ice-to-rock compositions, so this ratio is a reasonable assumption for the metal of gas giants formed outside an ice line. That said, at the extreme pressures found in gas giants, the difference in density between these various materials isn’t very large, so this choice is a second-order effect on the observables (radius, luminosity). The offerings for metal EOSs are more sparse than for H/He, but options include ANEOS (Thompson 1990), SESAME (Lyon and Johnson 1992), H2O-REOS (Nettelmann et al. 2010), Mazevet et al. (2019), and AQUA (Haldemann et al. 2020). MAGRATHEA (Huang et al. 2022) also provides a fair number of analytic equations of state for metals and a table-output function.
Multilayer Planets The discussion of the previous section has implicitly assumed that the planet in question is of uniform composition and convective throughout, but this is not necessarily the case. A major issue with this is that the core accretion model of planet formation requires a heavy-element core to form first in order to reach runaway gas accretion (Pollack et al. 1996). This has led to the use of a two layer model consisting of an isothermal metallic core and a convective H/He envelope (e.g., Marley et al. 2007; Lopez and Fortney 2014; Piso et al. 2015), especially in the sub-Neptune regime and for studies focused on the gas-accretion process (e.g., Lee et al. 2014). However, studies of the solar system planets have favored additional layers.
For Uranus and Neptune, these require at least the presence of an additional water-rich envelope layer between the core and outer envelope (see Fortney et al. 2010; Bailey and Stevenson 2021). Jupiter and Saturn are both believed to feature “fuzzy” cores (Wahl et al. 2017; Debras and Chabrier 2019; Movshovitz et al. 2020; Mankovich and Fuller 2021), which is to say that the composition gradually transitions from H/He poor to rich, rather than via a sharp boundary. This gradual transition may feature a staircase of , in which convective layers are separated by thin non-convecting boundaries with steep temperature and composition gradients (Leconte and Chabrier 2012). Such a staircase appears to be the natural outcome of initial composition gradients (Moll et al. 2016; Helled and Stevenson 2017; Fuentes and Cumming 2020), and could have a significant impact on the thermal evolution if the number of steps is large enough (Chabrier and Baraffe 2007; Leconte and Chabrier 2012). On the other hand, the number of layers evolves with time as layers exchange material and heat and potentially merge (Vazan et al. 2015, 2016), with the two diffusive fluxes across the boundary being positively related to one another (Moll et al. 2017). As such, even if planets formed with composition gradients that formed a large number of layers in a semi-convective staircase, the number of layers may have declined significantly as the planet subsequently evolved, reducing their impact on the present-day thermal structure of the planet.
For hot Jupiters, we lack direct evidence of the internal structure of the planets and must rely on comparisons with the solar system and the observed radii of the hot Jupiter population. Whether metals are mixed into the envelope or condensed into a (homogeneous) core has a relatively modest effect on the radius of the planet (Thorngren et al. 2016). On the other hand, many layers in a semiconvective staircase could inflate the radius by trapping heat in the planet’s deep interior (Leconte and Chabrier 2012). This would be a poor solution to the radius inflation problem, as it would produce the same inflation in cooler giant planets, yet this is not observed (Miller and Fortney 2011). Thus it is arguably most reasonable to assume the presence of a core sheathed in double-diffusive convective layers, but only a small number of layers in the outer envelope (as seen by Vazan et al. 2016, 2018, for Jupiter and Saturn). Such an arrangement would result in a warmer core than a two-layer model, but because heavy elements are much less thermally expansive than H/He, the effect on the planet’s radius would be somewhat limited.
3.2 Evolution Models
For planets consisting entirely of adiabatic and/or isothermal layers with a continuous temperature profile, the thermal state of the planet is reduced to a single variable. Typically this is the specific entropy of the envelope . A planet cools by releasing heat from the interior through its atmosphere, which depends on the thickness of the atmosphere (and therefore the mean molecular weight and gravity), its composition, and the incident flux from the star – highly irradiated atmospheres release heat from the interior more slowly. This can be specified as a total luminosity or an “intrinsic temperature” , which are related via , where is the Stefan-Boltzmann constant and R is the planet radius, usually at the radiative-convective boundary. Calculating the rate at which heat is released requires an atmosphere model that allows for both convective and radiative heat transport within the atmosphere down to the (RCB). This procedure is outside the scope of this chapter, but readers are instead directed to (Fortney et al. 2007) and the chapter of this book concerning radiative transfer. The rate of change of total energy in the planet can be written as
where is the hot Jupiter anomalous heating power (see Sec. 6.1 for estimates), is the heat generated by in the planet, is the release of tidal energy within the planet. This tidal heating could occur due to the damping of orbital eccentricity or obliquity between the planets rotational and orbital axes (Mardling 2007; Jackson et al. 2008, e.g.,); however, while it is likely important in certain cases (Millholland et al. 2020) it is not thought to explain the hot Jupiter radius anomaly on its own (Leconte et al. 2010), due to being short-lived on hot Jupiter orbits without an additional planet exciting eccentricity in the long term.
Heating from radioactive decays is commonly omitted from giant exoplanet studies, but work has been done to estimate its magnitude. Nettelmann et al. (2011) and Lopez et al. (2012) discuss its inclusion in their sub-Neptune models, considering four major radioactive isotopes: 235U, 238U, 232Th, and 40K, whose half-lives are 0.7, 4.46, 14, and 1.25 Gyr, respectively. The power generated over time is the product of the decay rate, decay energy, and the abundance of the isotope within the planet. The first two are known very precisely, but the latter is difficult to estimate. One option is to scale the isotopes with the modeled rock content in the same proportion as the protosolar abundances (from e.g., Anders and Grevesse 1989). Alternatively, Wang et al. (2020) suggest using the spectroscopic europium abundance of the host star as a proxy for 238U, 235U, and 232Th, since they are all refractory elements produced during r-process nucleosynthesis. Carrying the former analysis through, however, Thorngren et al. (2023) found that the contribution of a hot Jupiter’s luminosity from radioactive decays is extremely small and can be safely neglected.
Once a static structure model and the rate of change of energy in the interior has been calculated, calculating the change in envelope entropy with energy is a relatively simpler matter, requiring that one integrate the definition of temperature throughout the convective regions of the planet. If the planet contains a core, then for that region a specific heat should be used and the temperature change determined by the change in temperature of the envelope at the core boundary with s.
Here, is the specific heat of the core, is the core mass, is the core temperature (as in this case we have assumed it is isothermal), is the total mass of the planet, indicates mass contained within a shell at the given point in the planet, and is the temperature as a function of pressure and specific entropy . With Eqs. 7 and 8, one can calculate the rate of change of the specific entropy as the product . This may be integrated forward using a standard ODE integrator to obtain the of the planet. It’s worth noting that for hot Jupiters these equations can be somewhat “stiff” (an ODE term outside the scope of this review), so it can be helpful to select an integrator that tolerates stiff ODEs such as backward differentiation formula (BDF) integrators.
4 Heating Depth and Reinflation
An important factor in how gaseous planets respond to heating is the depth at which the heat is deposited. This is because convection can only move heat upwards – hot fluid is less dense and rises through a colder medium, assuming a uniform composition. As such, heat injected near the planet’s surface can only slow cooling, not reverse it (e.g., Wu and Lithwick 2013, for the Ohmic dissipation case). Komacek and Youdin (2017) goes further, showing that heat must be injected at least at the (RCB) in order to even slow the cooling; heat deposited above this point will be re-radiated out too quickly to affect the planet’s radius or thermal evolution significantly. While originally the RCB was thought to be very deep in the planet (Guillot et al. 1996), accounting for the planets’ hot interiors indicates that the RCB will be at around 10 bar for most hot Jupiters (Thorngren et al. 2019a), moving to lower pressures with increasing metallicity and/or decreasing surface gravity.
A prediction of any model based on heating near the surface of the planet (see Sec. 5) is therefore that planets will cool very slowly, but cannot grow larger ( ) if the flux is later increased (e.g., Wu and Lithwick 2013; Komacek et al. 2020). Lopez and Fortney (2016) observes that this is testable by examining whether giant planets around evolved stars are disproportionately larger as a result of the increase in stellar luminosity. Grunblatt et al. (2016, 2017) finds two such giant planets and observes that they are slightly larger than typical hot Jupiters would be without reinflation, though it would be helpful to find additional cases to verify the trend and potentially measure the rate at which reinflation occurs.
Post-main-sequence stars are not the only place to observe reinflation, however. Hartman et al. (2016) observed a relationship between hot Jupiter radii and the fractional age of their parents stars – the age relative to the main-sequence lifetime. This could result, he argued, from the gradual brightening of the stars on the main sequence. Thorngren et al. (2021) concur, observing that the degree of main-sequence brightening could indeed produce detectable increases in radius, and that these were consistent with the observed population. However, they were not able to estimate the rate that reinflation occurred at (which would be related to the deep-heating power), as main-sequence stellar brightening was too gradual. For this, the rapid brightening of post-main-sequence stars remains the best place to look.
5 Proposed Inflation Mechanisms
One of the earliest proposed explanations for the inflated radii of hot Jupiters was that the model atmosphere was significantly lower than in real hot Jupiters (Burrows et al. 2007, 2008). Such an error would cause models to cool too quickly, resulting in model planets that were too small. The advantages of this explanation include its lack of need for any new physics or a source of deep heating. However, it has a difficult time explaining the extremely large radii (up to ) that were subsequently discovered. As such, although this may be an important source of theoretical uncertainty in hot Jupiter atmosphere models, it does not appear to be the primary driver of their anomalous radii.
Another solution suggested early on was that of (Bodenheimer et al. 2001, 2003). Here, tides raised on a planet by the parent star would damp out the planet’s eccentricity and spin-orbit obliquity, depositing the energy reduction into the planet’s interior as heat. However, this is a finite reservoir of energy and the tidal timescales for hot Jupiters are in general quite short (Jackson et al. 2008), so these processes would only be important in the earliest stages of a hot Jupiter’s life. Additionally, tidal forces scale strongly with semimajor axis and radius, not the incident flux as has been observed. As such, while this may be important for some planets (Millholland and Spalding 2020) (not just hot Jupiters), it is not favored as a general solution (Leconte et al. 2010; Thorngren and Fortney 2018). On the other hand, Arras and Socrates (2009) suggest that tidal heating may not arise from eccentricity or obliquity, but rather the tidal pull on the planet’s thermal bulge (Socrates 2013) which is offset from the sub-stellar point by winds. In this model, energy is gradually pulled from the (enormous) reservoir of the planet’s orbit and fed into the deep interior. A key benefit of this model is that the deep heating could potentially reinflate the planet. However, follow-up studies differ as to whether the scaling is able to match the correlation between radius and flux (Sarkis et al. 2020; Thorngren and Fortney 2018).
Several proposed explanations for radius inflation can be broadly categorized together as fluid dynamic in nature. These share the common theme of arguing that the existing 1D atmospheric models are missing important physics that change the temperature-pressure profile of the deep atmosphere and suppress the release of heat from the interior. One such model is the Youdin and Mitchell (2010) proposal, in which the high-speed winds of the upper atmosphere generate turbulence beneath, dragging hot gas downward against the buoyancy force, producing a hotter and deeper RCB. Such an effect would slow the cooling process of the hot Jupiter substantially, though it would not by itself be capable of reinflating planets that had already cooled.
More recently, Tremblin et al. (2017) proposed that the inflation effect may be resolved in higher-dimensional atmosphere models. These 3D global circulation models (GCMs) are computationally expensive, especially when modeling the deep atmosphere where thermal timescales can be very long. Instead, they solve for the steady-state solutions in 2D at various latitudes. They find that longitudinally around the planet (steady-state flows, rather than turbulence) has the effect of burying heat deeper in the atmosphere than is indicated by 1D models. The result is suppressed cooling, hotter interiors, and therefore substantially larger radii. In fact, their model predictions were somewhat larger than the observed radii – they suggest among other things that this may be the result of the 2D model that will not occur in a full 3D model. Sainsbury-Martinez et al. (2019) studied this effect in 3D and supported their findings, suggesting that future studies use substantially longer integration times and initialize their simulations with temperature-pressure profiles above equilibrium for faster convergence.
A somewhat related proposal to the fluid dynamical models relies on a common feature of GCM results for hot Jupiters: a fast eastward (super-rotating) jet stream (e.g., Showman and Polvani 2011) carrying gas around the tidally locked planet. In the presence of a planetary magnetic field, this jet may cause ions (especially Na) and free electrons to drift such that a current flows through the atmosphere from the polar regions to the equator (Batygin and Stevenson 2010). This current flows back to the equator through the deeper parts of the atmosphere, depositing energy as a result of the material’s electrical resistivity. This process, called , has been proposed as a source of the heating required to inflate hot Jupiters (Batygin and Stevenson 2010; Batygin et al. 2011).
Originally it was thought that this would result in ultra-hot Jupiters inflating to the point that they overflow their Roche lobes, but subsequent studies disagree (Spiegel and Burrows 2013); Menou (2012); Rauscher and Menou (2013) found that magnetic braking can significantly reduce the velocity of the atmospheric winds, in turn reducing the effectiveness of Ohmic dissipation. The peak efficiency of turning incident light energy into interior heating was found to occur at K by Menou (2012) and K by Rogers and Komacek (2014). Additionally, it is worth noting that this heating mechanism is fairly shallow, perhaps around 10 bars (Rogers and Komacek 2014), so it is not capable of causing reinflation (Wu and Lithwick 2013, see Sec. 4). Pu and Valencia (2017) investigated the process for hot Neptunes and found that Ohmic dissipation may operate in a similar manner in those objects as in hot Jupiters. On the other hand, Huang and Cumming (2012) modeled the heating from Ohmic dissipation and instead argue that the total heating is not sufficient to explain the radii observed.
6 Observable Effects
In this section I’ll review some of the observation-based constraints that have been made to the nature of the hot Jupiter radius anomaly. Because of the extremely large number of diverse observations made of hot Jupiters, I’ll leave to other chapters observations which do not directly relate to the inflationary mechanism, especially transmission and emission spectra (discussed elsewhere in this book).
6.1 Population Trends in Radius
Given the large and growing population of observed hot Jupiters, a helpful approach to understanding them has been to consider their properties as a population, averaging out individual variations in metallicity or observational errors. Laughlin et al. (2011) did this by calculating the radius anomaly , the radius minus the expected radius for an evolved metal-free planet, for the 90 transiting hot Jupiters known at the time. They found that the anomaly was strongly related to the effective temperature, with , and argue that this appears consistent with the Ohmic dissipation model. While he supports the observation of a correlation between radius anomaly and incident flux, a key result is that the heating is not a fixed fraction of the incident flux, but one that itself increases with flux.
Following a similar thread, Thorngren and Fortney (2018) analyzed the then 318 hot Jupiters with masses and radii, directly including the internal heating power as a structure model parameter. The degeneracy between heating and composition was resolved by imposing the Thorngren et al. (2016) mass-metallicity relation as a prior. For individual planets, this doesn’t resolve the degeneracy fully, but applied to a large population it becomes possible to measure heating trends. They found that the as a fraction of flux rose to a peak of about 2.5% at 1570 K ( erg s-1 cm-1) and then declines to below 1% for the ultra-hot ( K) planets. They found a good fit using a Gaussian in log-flux to predict the heating:
Here, the amplitude is , the center is at , and the width is . Interestingly, there are hints of this decline in heating efficiency for the ultra-hot Jupiters in Laughlin et al. (2011) (that paper’s Fig. 2.), but there were too few such planets known at the time for this to have been statistically significant.
Sarkis et al. (2020) conducted a similar study, seeking to extract the heating efficiency from the population of hot Jupiters with a Bayesian hierarchical model. Unlike Thorngren and Fortney (2018), they included a number of rigorous checks including whether they could retrieve the modeled value from a synthesized population (they did), checks for prior sensitivity (their results were insensitive), and more extensive comparisons with predicted heating from various models. They too found a peak in heating efficiency of about 2.5%, but at a higher temperature of K; this difference does not appear to be due to choice of prior or the statistical framework, but may be a difference in atmosphere models. They present several parameterizations for estimating , but we’ll just reproduce the Gaussian here (in terms of ) for reference and comparison:
where is the equilibrium temperature, is the peak heating efficiency, is the temperature where the peak is located, and is the width of the Gaussian shape. Using a linear prior on (they also tried using a log prior), they found , , and . Comparing their inferred heating with model predictions, they agreed that the Ohmic heating found by Huang and Cumming (2012) was not sufficient to explain the radii, but that thermal tides (Socrates 2013) and temperature advection Tremblin et al. (2017) potentially could. That said, they did still find that the latter two, especially thermal tides, over-predicted the heating for ultra-hot Jupiters.
For the assumptions made by both Thorngren and Fortney (2018) and Sarkis et al. (2020) (deep heating), the planet rapidly approaches an equilibrium state in which the intrinsic luminosity of the planet is equal to the anomalous heating. This implies a direct relationship with the intrinsic temperature of the planet:
where is the Stefan-Boltzmann constant. Remember that depends on flux (or ), so this is not to be understood as a linear relationship between and . Still, using one of the formulas for allows one to calculate the relatively easily (e.g., Thorngren and Fortney 2019), which can be particularly helpful for studies where an estimate of it is needed; we’ll discuss such applications in the next section.
6.2 Intrinsic Temperature and Radiative-Convective Boundary
A hot interior has the potential to substantially affect the structure and heat flow through the planet’s atmosphere. The (see also Sec. 3.2) is a way to describe the energy entering the atmosphere from below. It participates in the heat balance of the atmosphere as:
That is, the heat exiting the planet is the sum of the incoming light energy plus the heat from the interior . Astronomers often rewrite this in terms of temperatures (Eq. 13) due to the popularity of the equilibrium temperature and effective temperatures. Note that is the intrinsic temperature and not the interior temperature, which varies with pressure! Nowhere in the interior will the actual temperature be as low as the intrinsic temperature. For cool giant planets, is initially large and drops over time as the planet cools; however, for hot Jupiters the anomalous heating may replace some or all of the heat going out, resulting in a steady-state intrinsic temperature in the hundreds of kelvin (Thorngren and Fortney 2018; Sarkis et al. 2020). On the other hand, if the large radii are caused by delayed cooling the intrinsic temperature would be by definition smaller than that predicted by existing atmosphere models.
The intrinsic temperature can have detectable impacts on the observable atmosphere (see Fortney et al. 2020; Gao et al. 2020); for example Sing et al. (2019a) argue that the presence of Fe and Mg in the atmosphere of WASP-121 (which they observed with HST) requires a of at least 500 K to avoid the atoms condensing in the atmosphere too deep to be observed. This is consistent with the deep heating case – Sarkis et al. (2020) calculate a K for that planet. The presence of absence of these cold traps is therefore a useful diagnostic of the intrinsic temperature of some planets (Parmentier et al. 2016; Fortney et al. 2020). The atmospheric chemistry may also be affected by , as higher values result in a higher , pushing the chemical abundances further from equilibrium. The intrinsic temperature is also an important input in global circulation models (GCMs) of giant planets (May et al. 2021; Steinrueck et al. 2021; Lee et al. 2022, e.g.,), though in that field it is referred to as internal temperature.
The intrinsic temperature and equilibrium temperature also directly affect the location of the (RCB, see Sec. 4): a higher pushes the RCB to higher pressures, whereas a high pushes it to lower pressures (e.g., Guillot et al. 1996; Fortney et al. 2007; Thorngren and Fortney 2019). When hot Jupiters were initially discovered, it was believed that their very high equilibrium temperatures would move the RCB very deep into the planet, at around 1 kilobar (e.g., Guillot et al. 1996). However, these models were based on RV discoveries, which had not yet revealed the radius anomaly problem. If large radii indeed imply large intrinsic temperatures, Thorngren et al. (2019a) and Sarkis et al. (2020) both find that the RCB is pushed back up to 1-100 bars, depending on the planets’ exact equilibrium temperature and surface gravity. As Sarkis et al. (2020) note, heat must be deposited below the RCB to have a significant impact on the radius (Komacek et al. 2017), so a shallower RCB imposes less stringent requirements on the pressures of heat deposition. They argue further that the same argument could be applied to the temperature advection mechanism of Tremblin et al. (2017).
6.3 Magnetic Fields
The s of exoplanets are poorly understood due to the difficulty in making direct measurements of them, but is thought to relate to the rotation rate, intrinsic temperature, or both (Reiners and Christensen 2010). Hot Jupiters are expected to be tidally locked Jackson et al. (2008), so even on their short periods the rotation rates will be lower than those of Jupiter and Saturn. On the other hand, their intrinsic temperatures appear to be substantially higher (Sec. 6.2). Assuming that hot Jupiter radii are explained by injected heat, Yadav and Thorngren (2017) applied the Christensen et al. (2009) scaling relations to infer hot Jupiter magnetic field strengths. The mean field on the dynamo surface from Christensen et al. (2009) was:
where is the planet mass, is the planet’s intrinsic luminosity (use the formulas from Sec. 6.2 and ), and R is the planet’s radius. For hot Jupiters, this results in polar surface field strengths of roughly 50-200 G (see Fig. 5), depending strongly on the incident stellar flux and planet mass. Such fields would be substantially stronger than those seen in Jupiter or Saturn, and could have impacts on the atmospheric flows (e.g., Beltz et al. 2023). These results are a reasonable estimate for what hot Jupiter fields might be, but they could be improved on by considering how the rotation rates factor in (e.g., Jeffries et al. 2011, for stars) and a more careful treatment of where the dynamo surface is.
An important angle for observationally constraining hot Jupiter magnetic fields is the interactions between the fields of the planet and its parent star. The is known to be a tracer of magnetic activity in stellar chromospheres (e.g., Schrijver et al. 1989), and so it has been hypothesized as a potential avenue for measuring star-planet magnetic interactions (see Lanza 2008; Poppenhaeger 2019), though the point of peak activity may not be directly facing the planet. Subsequent observations have found cases where the Ca II K line varies at the orbital period of the planet (e.g., Shkolnik et al. 2003, 2005), supporting the hypothesis that it is driven by star-planet interactions. An important development was made by Cauley et al. (2019), who calculated the magnetic field strengths implied by the Ca II K line modulation amplitudes for four hot Jupiters and found strong magnetic fields with surface strengths of 20-120 G, consistent (if possibly slightly lower) with the intrinsic temperature-based scaling laws used by Yadav and Thorngren (2017). Thus it provides observational support for the hypothesis that hot Jupiter magnetic fields are significantly stronger than those of cooler giants like Jupiter and Saturn.
The magnetic field of the planet is also thought to have an effect on the mass loss processes observed on some hot Jupiters (see Dos Santos 2021, for an overview). Adams (2011) show how magnetic fields can inhibit the rate of XUV-driven mass loss from giant planets with even relatively modest magnetic fields ( G); Owen and Adams (2014) agree and further argue that planetary magnetism will largely prevent mass loss from the night-side by inhibiting flows from the day-side (where the XUV irradiation is absorbed) around the planet. Additionally, the resulting outflows will be sculpted by the planetary magnetic fields; most importantly, escaping plasma follows the magnetic fields lines until it reaches the planet’s Alfvén surface, preventing mass from escaping from the regions away from the poles where the field lines never cross the Alfvén surface (Adams 2011; Carolan et al. 2021). Using these patterns, Schreyer et al. (2023) argues that the blue-shifting observed from escaping gas is tentative evidence for a magnetic field weak enough ( G) that gas can escape from the night side, as day-side escape would be red-shifted. The growing number of observations of gas escaping from hot Jupiters makes this a promising area for further theoretical work as well.
The use of radio observations has been suggested as a more direct approach for measuring the magnetic fields of hot Jupiters (Zarka 2007; Weber et al. 2018) based on the successful observation of the aurorae of Jupiter in radio (see Zarka 1998). Such searches would be made easier if the planets had a moon as a source of ions to supply the aurorae (Narang et al. 2023), but such moons are more likely for young distant objects than for hot Jupiters, where they would not be stable on long timescales (Barnes and O’Brien 2002). Searches have been conducted mainly for young, widely-separated planets (e.g., Bastian et al. 2018; Shiohira et al. 2024; Turner et al. 2024). These searches have not yet identified signals clearly attributable to the planet, but efforts are ongoing and still promising due to the successes in observing such signals from brown dwarfs (e.g., Kao et al. 2018).
6.4 Tidal Effects
Hot Jupiters on particularly close orbits are likely to be tidally distorted into a non-spherical shape by their parent stars. Unlike cooler giant planets, whose rotation can warp them into oblate spheroids, hot Jupiters are pulled into mostly-prolate (e.g., Li et al. 2010) tri-axial ellipsoids, where each of the three axes (sub-stellar, polar, and prograde) has a different radius. Estimating the self-consistent equilibrium shape theoretically is numerically expensive, but has been done for some of the most tidally-affected hot Jupiters by Wahl et al. (2021) using the concentric MacLaurin spheroid method. For moderately distorted planets, a good and much faster approximation of the shape can instead be achieved using the Love number (Love 1909; Leconte et al. 2011; Akinsanmi et al. 2019; Hellard et al. 2019), which is obtained from a static structure model via a relatively simple ODE (e.g., Buhler et al. 2016):
where is an intermediate quantity, is the radius in the planet, is the total radius of the planet, is the density at radius , and is the average density within the sphere of radius . The ODE is to be integrated from the initial conditions radially outwards from , so that may be plugged into Eq. 16 to obtain . Determining the Love number for hot Jupiters is of particular interest because, as a measure of how concentrated the planet’s density is towards the center, it can break the degeneracy between composition and specific entropy without any assumptions about the evolutionary history (e.g., Hellard et al. 2019)!
Observation of the tidal distortion is challenging but possible, and several avenues are available. First, the shape affects the shape of the transit by extending the amount of time spent in ingress and egress and a subtle decrease in transit depth towards mid-transit as the long axis of the planet faces the viewer (Leconte et al. 2011). The latter effect is difficult to separate from the effects of limb darkening, but the ingress and egress duration should be measurable in favorable cases (Akinsanmi et al. 2019). Barros et al. (2022) have claimed a 3-sigma detection of the shape of WASP-103 b using this approach. An alternative approach is a careful treatment of the the phase curve of the planet, which is affected by the shape even more strongly than the transit is (Akinsanmi et al. 2023); however, this signal must be carefully disentangled from others, including the planet temperature vs latitude, stellar deformation by the planet, and Doppler beaming of the emitted light.
A case of particular interest is that of HAT-P-13 b and c. Using a secular dynamics analysis, (Batygin et al. 2009) showed that the system fairly quickly reaches a steady-state in which gravitational interactions between b and c transfer orbital energy into the eccentricity of b, which is then damped by circularization tides onto a constant value of . More importantly, the two orbits precess together, and while the details of the secular dynamics arguments are too involved to go into here, the result is that the eccentricity of b is directly related to it’s value of ! Follow-up by (Buhler et al. 2016) used HST secondary transits to determine an eccentricity of , implying , which corresponds to a most likely core mass of and a 1-sigma upper limit of , reasonable numbers for the object. This is so far the only constraint on an exoplanet’s precise enough to be useful for constraining its interior structure, but it seems likely that more will be measured, especially as more long-period companions of known exoplanets are discovered.
7 Conclusion
The hot Jupiter radius inflation problem continues to be one of the longest-standing open questions in exoplanet science; however, over the last two decades we have made significant progress in collecting observational evidence for its cause. To review the key points of evidence from previous sections:
1.
The entire population of hot Jupiters shows signs of inflation; the lower-radius examples are consistent with higher metallicity, not a lack of inflation (Laughlin et al. 2011).
2.
The degree of radius inflation is correlated with incident stellar flux, not period or semimajor axis (Weiss et al. 2013).
- 3.
- 4.
5.
The absence of cold traps on WASP-121 b indicates a high intrinsic temperature (Sing et al. 2019b).
- 6.
7.
Young hot Jupiters appear to be slightly larger than their older analogues (Thorngren et al. 2021).
Comparing these observations with the proposed inflation mechanisms in Sec. 5, it is evident that there is a tension between them. The presence of a uniform inflation effect strongly related to incident flux suggests that the inflation mechanism is closely tied to atmospheric dynamics, such as Ohmic dissipation or temperature advection. However, these mechanisms are not able to explain the reinflation of hot Jupiters seen in several studies. Though thermal tides models are able to bring about reinflation, the heating predicted by them does not appear to match the radii of ultra-hot Jupiters (Sarkis et al. 2020), seeming to overinflate them (Thorngren and Fortney 2018). One possibility is that we are missing an important effect that limits the impact of thermal tides at very high equilibrium temperatures – perhaps magnetic breaking prevents these objects’ hot spots from moving as far.
Alternatively, we may need to consider multiple mechanisms operating at the same time. If both thermal tides and Ohmic dissipation contribute significantly, for example, they might match the observations by having Ohmic dissipation regulate heat flow out of the planet while modest thermal tides allow the planet to reinflate. The key point here is that we shouldn’t require that any one inflation mechanism solve the entire problem on its own – they may be acting in concert with one another.
Let’s conclude with some promising areas of investigation. On the observational side, the discovery of more hot Jupiters is always helpful for doing population statistics, but those at very young ages or around post-main sequence stars will be particularly valuable as constraints on the radius evolution of hot Jupiters. For individual planet studies, any atmosphere observations that can measure or constrain the intrinsic temperature (as in Sing et al. 2019b) are of particular interest, as would constraints on the atmospheric metallicity, as these help to narrow down the range of plausible interiors. Observations of the magnetic field strengths of hot Jupiters is similarly helpful, be they through star-planet interactions, direct radio observation, or some other method. Finally, a sufficiently precise observation of the prolate shape of a hot Jupiter (or equivalently its Love number) would provide a valuable new angle of constraint on the deep interior.
For theoretical work, a particularly noteworthy need is a revisitation of the thermal tides model, which has not been substantially developed since Socrates (2013). Our improved understanding of hot Jupiter circulation, including the day-night contrasts and hot-spot offsets, warrant further work. Likewise, GCMs extending down into the deep atmosphere on long enough integration times for various masses and instellation would be computationally expensive but extremely valuable in addressing the temperature advection mode. Considering also the effect of magnetic fields worsens the computational expense, but may be necessary to understand how these revised atmospheric flows interact with Ohmic dissipation. Finally, the rapidly enlarging sample of observed hot Jupiters and recent developments in H/He equations of state make it worth recalculating the trends in heating with incident flux with particular focus on the cooling and reinflation rates.
Between the abundance of new observations and the ongoing theoretical work to understand them, this is a very exciting time for the study of hot Jupiters. The coming years will likely see further dramatic improvements in our understanding of these planets.
8 Cross-References
Radiative Transfer for Exoplanet Atmospheres - Heng & Marley
Exoplanet Atmosphere Measurements from Transmission Spectroscopy and Other Planet Star Combined Light Observations - Kreidberg
Tidal Star-Planet Interactions: A Stellar and Planetary Perspective - Mathis
Acknowledgements.
Thanks go to the Allan C. and Dorothy H. Davis Postdoctoral Fellowship at Johns Hopkins University for support, and to Stephen Schmidt for extensive proofreading and readability suggestions.References
- Adams (2011) Adams FC (2011) Magnetically Controlled Outflows from Hot Jupiters. The Astrophysical Journal 730(1):27
- Akinsanmi et al. (2019) Akinsanmi B, Barros SCC, Santos NC et al. (2019) Detectability of shape deformation in short-period exoplanets. Astronomy and Astrophysics 621:A117
- Akinsanmi et al. (2023) Akinsanmi B, Lendl M, Boue G Barros SCC (2023) On the effect of tidal deformation on planetary phase curves. DOI 10.48550/arXiv.2310.03553
- Alonso et al. (2004) Alonso R, Brown TM, Torres G et al. (2004) TrES-1: The Transiting Planet of a Bright K0 V Star. The Astrophysical Journal 613:L153–L156
- Anders and Grevesse (1989) Anders E Grevesse N (1989) Abundances of the elements: Meteoritic and solar. Geochimica et Cosmochimica Acta 53:197–214
- Arras and Socrates (2009) Arras P Socrates A (2009) Thermal Tides in Short Period Exoplanets. ArXiv e-prints 0901:arXiv:0901.0735
- Asplund et al. (2009) Asplund M, Grevesse N, Sauval AJ Scott P (2009) The Chemical Composition of the Sun. Annual Review of Astronomy and Astrophysics 47:481–522
- Auvergne et al. (2009) Auvergne M, Bodin P, Boisnard L et al. (2009) The CoRoT satellite in flight: Description and performance. Astronomy and Astrophysics 506:411–424
- Bailey and Stevenson (2021) Bailey E Stevenson DJ (2021) Thermodynamically Governed Interior Models of Uranus and Neptune. The Planetary Science Journal 2:64
- Bakos et al. (2004) Bakos G, Noyes RW, Kovács G et al. (2004) Wide-Field Millimagnitude Photometry with the HAT: A Tool for Extrasolar Planet Detection. Publications of the Astronomical Society of the Pacific 116:266–277
- Bakos et al. (2007) Bakos GÁ, Noyes RW, Kovács G et al. (2007) HAT-P-1b: A Large-Radius, Low-Density Exoplanet Transiting One Member of a Stellar Binary. The Astrophysical Journal 656:552–559
- Bakos et al. (2013) Bakos GÁ, Csubry Z, Penev K et al. (2013) HATSouth: A Global Network of Fully Automated Identical Wide-Field Telescopes1. Publications of the Astronomical Society of the Pacific 125(924):154–182
- Baraffe et al. (2003) Baraffe I, Chabrier G, Barman TS, Allard F Hauschildt PH (2003) Evolutionary models for cool brown dwarfs and extrasolar giant planets. The case of HD 209458. Astronomy and Astrophysics 402:701–712
- Barnes and O’Brien (2002) Barnes JW O’Brien DP (2002) Stability of Satellites around Close-in Extrasolar Giant Planets. The Astrophysical Journal 575:1087–1093
- Barros et al. (2022) Barros SCC, Akinsanmi B, Boué G et al. (2022) Detection of the tidal deformation of WASP-103b at 3 with CHEOPS. Astronomy and Astrophysics 657:A52
- Bastian et al. (2018) Bastian TS, Villadsen J, Maps A, Hallinan G Beasley AJ (2018) Radio Emission from the Exoplanetary System Eridani. The Astrophysical Journal 857(2):133
- Batygin and Stevenson (2010) Batygin K Stevenson DJ (2010) Inflating Hot Jupiters with Ohmic Dissipation. The Astrophysical Journal Letters 714(2):L238
- Batygin et al. (2009) Batygin K, Bodenheimer P Laughlin G (2009) Determination of the Interior Structure of Transiting Planets in Multiple-Planet Systems. The Astrophysical Journal Letters 704:L49–L53
- Batygin et al. (2011) Batygin K, Stevenson DJ Bodenheimer PH (2011) Evolution of Ohmically Heated Hot Jupiters. The Astrophysical Journal 738:1
- Beltz et al. (2023) Beltz H, Rauscher E, Kempton EMR, Malsky I Savel AB (2023) Magnetic Effects and 3D Structure in Theoretical High-resolution Transmission Spectra of Ultrahot Jupiters: The Case of WASP-76b. The Astronomical Journal 165(6):257
- Bodenheimer et al. (2000) Bodenheimer P, Hubickyj O Lissauer JJ (2000) Models of the in Situ Formation of Detected Extrasolar Giant Planets. Icarus 143:2–14
- Bodenheimer et al. (2001) Bodenheimer P, Lin DNC Mardling RA (2001) On the Tidal Inflation of Short-Period Extrasolar Planets. The Astrophysical Journal 548:466–472
- Bodenheimer et al. (2003) Bodenheimer P, Laughlin G Lin DNC (2003) On the Radii of Extrasolar Giant Planets. The Astrophysical Journal 592:555–563
- Borucki et al. (2010) Borucki WJ, Koch D, Basri G et al. (2010) Kepler Planet-Detection Mission: Introduction and First Results. Science 327:977
- Buhler et al. (2016) Buhler PB, Knutson HA, Batygin K et al. (2016) Dynamical Constraints on the Core Mass of Hot Jupiter HAT-P-13b. The Astrophysical Journal 821:26
- Burrows et al. (2003) Burrows A, Sudarsky D Hubbard WB (2003) A Theory for the Radius of the Transiting Giant Planet HD 209458b. The Astrophysical Journal 594:545–551
- Burrows et al. (2004) Burrows A, Hubeny I, Hubbard WB, Sudarsky D Fortney JJ (2004) Theoretical Radii of Transiting Giant Planets: The Case of OGLE-TR-56b. The Astrophysical Journal 610(1):L53
- Burrows et al. (2007) Burrows A, Hubeny I, Budaj J Hubbard WB (2007) Possible Solutions to the Radius Anomalies of Transiting Giant Planets. The Astrophysical Journal 661:502–514
- Burrows et al. (2008) Burrows A, Ibgui L Hubeny I (2008) Optical Albedo Theory of Strongly Irradiated Giant Planets: The Case of HD 209458b. The Astrophysical Journal 682(2):1277–1282
- Caldiroli et al. (2022) Caldiroli A, Haardt F, Gallo E et al. (2022) Irradiation-driven escape of primordial planetary atmospheres. II. Evaporation efficiency of sub-Neptunes through hot Jupiters. Astronomy and Astrophysics 663:A122
- Carolan et al. (2021) Carolan S, Vidotto AA, Hazra G, Villarreal D’Angelo C Kubyshkina D (2021) The effects of magnetic fields on observational signatures of atmospheric escape in exoplanets: Double tail structures. Monthly Notices of the Royal Astronomical Society 508(4):6001–6012
- Cauley et al. (2019) Cauley PW, Shkolnik EL, Llama J Lanza AF (2019) Magnetic field strengths of hot Jupiters from signals of star–planet interactions. Nature Astronomy p 1
- Chabrier and Baraffe (2007) Chabrier G Baraffe I (2007) Heat Transport in Giant (Exo)planets: A New Perspective. The Astrophysical Journal Letters 661:L81–L84
- Chabrier and Debras (2021) Chabrier G Debras F (2021) A New Equation of State for Dense Hydrogen-Helium Mixtures. II. Taking into Account Hydrogen-Helium Interactions. The Astrophysical Journal 917:4
- Chabrier et al. (1992) Chabrier G, Saumon D, Hubbard WB Lunine JI (1992) The molecular-metallic transition of hydrogen and the structure of Jupiter and Saturn. The Astrophysical Journal 391:817–826
- Chabrier et al. (2004) Chabrier G, Barman T, Baraffe I, Allard F Hauschildt PH (2004) The Evolution of Irradiated Planets: Application to Transits. The Astrophysical Journal 603(1):L53–L56
- Chabrier et al. (2019) Chabrier G, Mazevet S Soubiran F (2019) A New Equation of State for Dense Hydrogen–Helium Mixtures. The Astrophysical Journal 872:51
- Charbonneau et al. (2000) Charbonneau D, Brown TM, Latham DW Mayor M (2000) Detection of Planetary Transits Across a Sun-like Star. The Astrophysical Journal Letters 529:L45–L48
- Chen and Rogers (2016) Chen H Rogers LA (2016) Evolutionary Analysis of Gaseous Sub-Neptune-mass Planets with MESA. The Astrophysical Journal 831:180
- Chen and Kipping (2017) Chen J Kipping D (2017) Probabilistic Forecasting of the Masses and Radii of Other Worlds. The Astrophysical Journal 834(1):17
- Christensen et al. (2009) Christensen UR, Holzwarth V Reiners A (2009) Energy flux determines magnetic field strength of planets and stars. Nature 457:167–169
- Collier Cameron et al. (2007) Collier Cameron A, Bouchy F, Hébrard G et al. (2007) WASP-1b and WASP-2b: Two new transiting exoplanets detected with SuperWASP and SOPHIE. Monthly Notices of the Royal Astronomical Society 375:951–957
- Cumming et al. (2018) Cumming A, Helled R Venturini J (2018) The primordial entropy of Jupiter. Monthly Notices of the Royal Astronomical Society 477:4817–4823
- Debras and Chabrier (2019) Debras F Chabrier G (2019) New models of Jupiter in the context of Juno and Galileo. arXiv:190105697 [astro-ph]
- Demory and Seager (2011) Demory BO Seager S (2011) Lack of Inflated Radii for Kepler Giant Planet Candidates Receiving Modest Stellar Irradiation. The Astrophysical Journal Supplement Series 197(1):12
- Dos Santos (2021) Dos Santos LA (2021) Observations of planetary winds and outflows. Proceedings of the International Astronomical Union 17(S370):56–71
- Dwivedi et al. (2019) Dwivedi NK, Khodachenko ML, Shaikhislamov IF et al. (2019) Modelling atmospheric escape and MgII near-ultraviolet absorption of the highly irradiated hot Jupiter WASP-12b. Monthly Notices of the Royal Astronomical Society 487(3):4208–4220
- Ford and Rasio (2006) Ford EB Rasio FA (2006) On the Relation between Hot Jupiters and the Roche Limit. The Astrophysical Journal 638:L45–L48
- Fortney et al. (2007) Fortney JJ, Marley MS Barnes JW (2007) Planetary Radii across Five Orders of Magnitude in Mass and Stellar Insolation: Application to Transits. The Astrophysical Journal 659:1661–1672
- Fortney et al. (2010) Fortney JJ, Shabram M, Showman AP et al. (2010) Transmission Spectra of Three-Dimensional Hot Jupiter Model Atmospheres. The Astrophysical Journal 709:1396–1406
- Fortney et al. (2020) Fortney JJ, Visscher C, Marley MS et al. (2020) Beyond Equilibrium Temperature: How the Atmosphere/Interior Connection Affects the Onset of Methane, Ammonia, and Clouds in Warm Transiting Giant Planets. The Astronomical Journal 160:288
- Fuentes and Cumming (2020) Fuentes JR Cumming A (2020) Penetration of a cooling convective layer into a stably-stratified composition gradient: Entrainment at low Prandtl number. arXiv:200704265 [astro-ph, physics:physics]
- Gao et al. (2020) Gao P, Thorngren DP, Lee EKH et al. (2020) Aerosol composition of hot giant exoplanets dominated by silicates and hydrocarbon hazes. Nature Astronomy 4:951–956
- Gaudi et al. (2017) Gaudi BS, Stassun KG, Collins KA et al. (2017) A giant planet undergoing extreme-ultraviolet irradiation by its hot massive-star host. Nature 546(7659):514–518
- Grunblatt et al. (2016) Grunblatt SK, Huber D, Gaidos EJ et al. (2016) K2-97b: A (Re-?)Inflated Planet Orbiting a Red Giant Star. The Astronomical Journal 152:185
- Grunblatt et al. (2017) Grunblatt SK, Huber D, Gaidos E et al. (2017) Seeing Double with K2: Testing Re-inflation with Two Remarkably Similar Planets around Red Giant Branch Stars. The Astronomical Journal 154:254
- Guillot and Showman (2002) Guillot T Showman AP (2002) Evolution of “51 Pegasus b-like” planets. Astronomy and Astrophysics 385:156–165
- Guillot et al. (1996) Guillot T, Burrows A, Hubbard WB, Lunine JI Saumon D (1996) Giant Planets at Small Orbital Distances. The Astrophysical Journal 459:L35
- Haldemann et al. (2020) Haldemann J, Alibert Y, Mordasini C Benz W (2020) AQUA: A Collection of H$_2$O Equations of State for Planetary Models. Astronomy & Astrophysics 643:A105
- Hamer and Schlaufman (2019) Hamer JH Schlaufman KC (2019) Hot Jupiters are Destroyed by Tides While Their Host Stars are on the Main Sequence. arXiv e-prints p arXiv:1908.06998
- Hartman et al. (2016) Hartman JD, Bakos GÁ, Bhatti W et al. (2016) HAT-P-65b and HAT-P-66b: Two Transiting Inflated Hot Jupiters and Observational Evidence for the Reinflation of Close-in Giant Planets. The Astronomical Journal 152:182
- Hellard et al. (2019) Hellard H, Csizmadia S, Padovan S et al. (2019) Retrieval of the fluid Love number $k_2$ in exoplanetary transit curves. arXiv e-prints 1905:arXiv:1905.03,171
- Helled and Stevenson (2017) Helled R Stevenson D (2017) The Fuzziness of Giant Planets’ Cores. The Astrophysical Journal 840:L4
- Henry et al. (2000) Henry GW, Marcy GW, Butler RP Vogt SS (2000) A Transiting “51 Peg-like” Planet. The Astrophysical Journal Letters 529:L41–L44
- Hoeijmakers et al. (2018) Hoeijmakers HJ, Ehrenreich D, Heng K et al. (2018) Atomic iron and titanium in the atmosphere of the exoplanet KELT-9b. arXiv:180805653 [astro-ph, physics:physics]
- Howell et al. (2014) Howell SB, Sobeck C, Haas M et al. (2014) The K2 Mission: Characterization and Early Results. Publications of the Astronomical Society of the Pacific 126:398
- Huang et al. (2022) Huang C, Rice DR Steffen JH (2022) MAGRATHEA: An open-source spherical symmetric planet interior structure code. Monthly Notices of the Royal Astronomical Society 513:5256–5269
- Huang and Cumming (2012) Huang X Cumming A (2012) Ohmic Dissipation in the Interiors of Hot Jupiters. The Astrophysical Journal 757(1):47
- Ida and Lin (2004) Ida S Lin DNC (2004) Toward a Deterministic Model of Planetary Formation. I. A Desert in the Mass and Semimajor Axis Distributions of Extrasolar Planets. The Astrophysical Journal 604:388–413
- Jackson et al. (2008) Jackson B, Greenberg R Barnes R (2008) Tidal Heating of Extrasolar Planets. The Astrophysical Journal 681:1631–1638
- Jeffries et al. (2011) Jeffries RD, Jackson RJ, Briggs KR, Evans PA Pye JP (2011) Investigating coronal saturation and supersaturation in fast-rotating M-dwarf stars: Coronal supersaturation in M-dwarfs. Monthly Notices of the Royal Astronomical Society 411(3):2099–2112
- Johnson et al. (2010) Johnson JA, Aller KM, Howard AW Crepp JR (2010) Giant Planet Occurrence in the Stellar Mass-Metallicity Plane. Publications of the Astronomical Society of the Pacific 122:905
- Kao et al. (2018) Kao MM, Hallinan G, Pineda JS, Stevenson D Burgasser A (2018) The Strongest Magnetic Fields on the Coolest Brown Dwarfs. The Astrophysical Journal Supplement Series 237(2):25
- Komacek and Youdin (2017) Komacek TD Youdin AN (2017) Structure and Evolution of Internally Heated Hot Jupiters. The Astrophysical Journal 844:94
- Komacek et al. (2017) Komacek TD, Showman AP Tan X (2017) Atmospheric Circulation of Hot Jupiters: Dayside-Nightside Temperature Differences. II. Comparison with Observations. The Astrophysical Journal 835(2):198
- Komacek et al. (2020) Komacek TD, Thorngren DP, Lopez ED Ginzburg S (2020) Reinflation of Warm and Hot Jupiters. The Astrophysical Journal 893:36
- Konacki et al. (2003) Konacki M, Torres G, Jha S Sasselov DD (2003) An extrasolar planet that transits the disk of its parent star. Nature 421:507–509
- Kurokawa and Nakamoto (2014) Kurokawa H Nakamoto T (2014) Mass-loss Evolution of Close-in Exoplanets: Evaporation of Hot Jupiters and the Effect on Population. The Astrophysical Journal 783:54
- Lanza (2008) Lanza AF (2008) Hot Jupiters and stellar magnetic activity. Astronomy & Astrophysics 487(3):1163–1170
- Laughlin et al. (2011) Laughlin G, Crismani M Adams FC (2011) On the Anomalous Radii of the Transiting Extrasolar Planets. The Astrophysical Journal Letters 729:L7
- Leconte and Chabrier (2012) Leconte J Chabrier G (2012) A new vision of giant planet interiors: Impact of double diffusive convection. Astronomy & Astrophysics 540:A20
- Leconte et al. (2010) Leconte J, Chabrier G, Baraffe I Levrard B (2010) Is tidal heating sufficient to explain bloated exoplanets? Consistent calculations accounting for finite initial eccentricity. Astronomy & Astrophysics 516:A64
- Leconte et al. (2011) Leconte J, Chabrier G, Baraffe I Levrard B (2011) The radius anomaly in the planet/brown dwarf overlapping mass regime. In: European Physical Journal Web of Conferences, eprint: arXiv:1011.0336, vol 11, p 03004, DOI 10.1051/epjconf/20101103004
- Lee et al. (2014) Lee EJ, Chiang E Ormel CW (2014) Make Super-Earths, Not Jupiters: Accreting Nebular Gas onto Solid Cores at 0.1 AU and Beyond. The Astrophysical Journal 797:95
- Lee et al. (2022) Lee EJ, Fuentes JR Hopkins PF (2022) Establishing Dust Rings and Forming Planets within Them. The Astrophysical Journal 937:95
- Li et al. (2010) Li SL, Miller N, Lin DNC Fortney JJ (2010) WASP-12b as a prolate, inflated and disrupting planet from tidal dissipation. Nature 463:1054–1056
- Lopez and Fortney (2014) Lopez ED Fortney JJ (2014) Understanding the Mass-Radius Relation for Sub-neptunes: Radius as a Proxy for Composition. The Astrophysical Journal 792:1
- Lopez and Fortney (2016) Lopez ED Fortney JJ (2016) Re-inflated Warm Jupiters around Red Giants. The Astrophysical Journal 818:4
- Lopez et al. (2012) Lopez ED, Fortney JJ Miller N (2012) How Thermal Evolution and Mass-loss Sculpt Populations of Super-Earths and Sub-Neptunes: Application to the Kepler-11 System and Beyond. The Astrophysical Journal 761:59
- Love (1909) Love AEH (1909) The yielding of the earth to disturbing forces. Proceedings of the Royal Society of London Series A, Containing Papers of a Mathematical and Physical Character 82(551):73–88
- Lyon and Johnson (1992) Lyon SP Johnson JD (1992) SESAME: The Los Alamos National Laboratory equation of state database. Tech. Rep. LA-UR-92-3407, Los Alamos National Laboratory
- Mankovich and Fuller (2021) Mankovich C Fuller J (2021) A diffuse core in Saturn revealed by ring seismology. arXiv:210413385 [astro-ph]
- Mankovich et al. (2016) Mankovich C, Fortney JJ Moore KL (2016) Bayesian Evolution Models for Jupiter with Helium Rain and Double-diffusive Convection. The Astrophysical Journal 832:113
- Mardling (2007) Mardling RA (2007) Long-term tidal evolution of short-period planets with companions. Monthly Notices of the Royal Astronomical Society 382:1768–1790
- Marley et al. (2007) Marley MS, Fortney JJ, Hubickyj O, Bodenheimer P Lissauer JJ (2007) On the Luminosity of Young Jupiters. The Astrophysical Journal 655(1):541
- Marzari and Weidenschilling (2002) Marzari F Weidenschilling SJ (2002) Eccentric Extrasolar Planets: The Jumping Jupiter Model. Icarus 156:570–579
- Masset and Papaloizou (2003) Masset FS Papaloizou JCB (2003) Runaway Migration and the Formation of Hot Jupiters. The Astrophysical Journal 588:494–508
- May et al. (2021) May EM, Komacek TD, Stevenson KB et al. (2021) Spitzer Phase-curve Observations and Circulation Models of the Inflated Ultrahot Jupiter WASP-76b. The Astronomical Journal 162(4):158
- Mayor and Queloz (1995) Mayor M Queloz D (1995) A Jupiter-mass companion to a solar-type star. Nature 378:355–359
- Mazevet et al. (2019) Mazevet S, Licari A, Chabrier G Potekhin AY (2019) Ab initio based equation of state of dense water for planetary and exoplanetary modeling. Astronomy and Astrophysics 621:A128
- Menou (2012) Menou K (2012) Magnetic Scaling Laws for the Atmospheres of Hot Giant Exoplanets. The Astrophysical Journal 745:138
- Militzer and Hubbard (2013) Militzer B Hubbard WB (2013) Ab Initio Equation of State for Hydrogen-Helium Mixtures with Recalibration of the Giant-planet Mass-Radius Relation. The Astrophysical Journal 774:148
- Miller and Fortney (2011) Miller N Fortney JJ (2011) The Heavy-element Masses of Extrasolar Giant Planets, Revealed. The Astrophysical Journal Letters 736:L29
- Millholland et al. (2020) Millholland S, Petigura E Batygin K (2020) Tidal Inflation Reconciles Low-density Sub-Saturns with Core Accretion. The Astrophysical Journal 897:7
- Millholland and Spalding (2020) Millholland SC Spalding C (2020) Formation of Ultra-short-period Planets by Obliquity-driven Tidal Runaway. The Astrophysical Journal 905:71
- Moll et al. (2016) Moll R, Garaud P Stellmach S (2016) A New Model for Mixing by Double-diffusive Convection (Semi-convection). III. Thermal and Compositional Transport through Non-layered ODDC. The Astrophysical Journal 823:33
- Moll et al. (2017) Moll R, Garaud P, Mankovich C Fortney JJ (2017) Double-diffusive Erosion of the Core of Jupiter. The Astrophysical Journal 849:24
- Movshovitz et al. (2020) Movshovitz N, Fortney JJ, Mankovich C, Thorngren D Helled R (2020) Saturn’s Probable Interior: An Exploration of Saturn’s Potential Interior Density Structures. The Astrophysical Journal 891:109
- Murray-Clay et al. (2009) Murray-Clay RA, Chiang EI Murray N (2009) Atmospheric Escape From Hot Jupiters. The Astrophysical Journal 693:23–42
- Naef et al. (2001) Naef D, Latham DW, Mayor M et al. (2001) HD 80606 b, a planet on an extremely elongated orbit. Astronomy & Astrophysics 375(2):L27–L30
- Narang et al. (2023) Narang M, Oza AV, Hakim K et al. (2023) Radio-loud Exoplanet-exomoon Survey: GMRT Search for Electron Cyclotron Maser Emission. The Astronomical Journal 165(1):1
- Nettelmann et al. (2008) Nettelmann N, Holst B, Kietzmann A et al. (2008) Ab Initio Equation of State Data for Hydrogen, Helium, and Water and the Internal Structure of Jupiter. The Astrophysical Journal 683:1217–1228
- Nettelmann et al. (2010) Nettelmann N, Kramm U, Redmer R Neuhäuser R (2010) Interior structure models of GJ 436b. Astronomy and Astrophysics 523:A26
- Nettelmann et al. (2011) Nettelmann N, Fortney JJ, Kramm U Redmer R (2011) Thermal Evolution and Structure Models of the Transiting Super-Earth GJ 1214b. The Astrophysical Journal 733:2
- Newton et al. (2019) Newton ER, Mann AW, Tofflemire BM et al. (2019) TESS Hunt for Young and Maturing Exoplanets (THYME): A Planet in the 45 Myr Tucana-Horologium Association. The Astrophysical Journal 880:L17
- Owen and Adams (2014) Owen JE Adams FC (2014) Magnetically controlled mass-loss from extrasolar planets in close orbits. Monthly Notices of the Royal Astronomical Society 444:3761–3779
- Parmentier et al. (2016) Parmentier V, Fortney JJ, Showman AP, Morley C Marley MS (2016) Transitions in the Cloud Composition of Hot Jupiters. The Astrophysical Journal 828(1):22
- Paxton et al. (2013) Paxton B, Cantiello M, Arras P et al. (2013) Modules for Experiments in Stellar Astrophysics (MESA): Planets, Oscillations, Rotation, and Massive Stars. The Astrophysical Journal Supplement Series 208(1):4
- Piso et al. (2015) Piso AMA, Youdin AN Murray-Clay RA (2015) Minimum Core Masses for Giant Planet Formation With Realistic Equations of State and Opacities. The Astrophysical Journal 800(2):82
- Pollacco et al. (2006) Pollacco DL, Skillen I, Collier Cameron A et al. (2006) The WASP Project and the SuperWASP Cameras. Publications of the Astronomical Society of the Pacific 118:1407–1418
- Pollack et al. (1996) Pollack JB, Hubickyj O, Bodenheimer P et al. (1996) Formation of the Giant Planets by Concurrent Accretion of Solids and Gas. Icarus 124(1):62–85
- Poppenhaeger (2019) Poppenhaeger K (2019) How stars and planets interact: A look through the high-energy window. Astronomische Nachrichten 340(4):329–333
- Pu and Valencia (2017) Pu B Valencia D (2017) Ohmic Dissipation in Mini-Neptunes. The Astrophysical Journal 846(1):47
- Rauscher and Menou (2013) Rauscher E Menou K (2013) Three-Dimensional Atmospheric Circulation Models of HD 189733b and HD 209458b with Consistent Magnetic Drag and Ohmic Dissipation. The Astrophysical Journal 764(1):103
- Reiners and Christensen (2010) Reiners A Christensen UR (2010) A magnetic field evolution scenario for brown dwarfs and giant planets. Astronomy & Astrophysics 522:A13
- Ricker et al. (2014) Ricker GR, Winn JN, Vanderspek R et al. (2014) Transiting Exoplanet Survey Satellite. Journal of Astronomical Telescopes, Instruments, and Systems 1(1):014,003
- Rogers and Komacek (2014) Rogers TM Komacek TD (2014) Magnetic Effects in Hot Jupiter Atmospheres. The Astrophysical Journal 794:132
- Sainsbury-Martinez et al. (2019) Sainsbury-Martinez F, Wang P, Fromang S et al. (2019) Idealised simulations of the deep atmosphere of hot Jupiters: Deep, hot adiabats as a robust solution to the radius inflation problem. Astronomy & Astrophysics 632:A114
- Sarkis et al. (2020) Sarkis P, Mordasini C, Henning T, Marleau GD Mollière P (2020) Evidence of Three Mechanisms Explaining the Radius Anomaly of Hot Jupiters. arXiv e-prints 2009:arXiv:2009.04,291
- Schreyer et al. (2023) Schreyer E, Owen JE, Spake JJ, Bahroloom Z Di Giampasquale S (2023) Using helium 10 830 Å transits to constrain planetary magnetic fields. Monthly Notices of the Royal Astronomical Society 527(3):5117–5130
- Schrijver et al. (1989) Schrijver CJ, Cote J, Zwaan C Saar SH (1989) Relations between the photospheric magnetic field and the emission from the outer atmospheres of cool stars. I - The solar CA II K line core emission. The Astrophysical Journal 337:964
- Serenelli and Basu (2010) Serenelli AM Basu S (2010) Determining the Initial Helium Abundance of the Sun. The Astrophysical Journal 719(1):865
- Shiohira et al. (2024) Shiohira Y, Fujii Y, Kita H et al. (2024) A search for auroral radio emission from Pictoris b. Monthly Notices of the Royal Astronomical Society 528(2):2136–2144
- Shkolnik et al. (2003) Shkolnik E, Walker GAH Bohlender DA (2003) Evidence for Planet-induced Chromospheric Activity on HD 179949. The Astrophysical Journal 597(2):1092–1096
- Shkolnik et al. (2005) Shkolnik E, Walker GAH, Bohlender DA, Gu PG Kurster M (2005) Hot Jupiters and Hot Spots: The Short- and Long-Term Chromospheric Activity on Stars with Giant Planets. The Astrophysical Journal 622(2):1075–1090
- Showman and Polvani (2011) Showman AP Polvani LM (2011) Equatorial Superrotation on Tidally Locked Exoplanets. The Astrophysical Journal 738(1):71
- Sing et al. (2019a) Sing DK, Lavvas P, Ballester GE et al. (2019a) The Hubble Space Telescope PanCET Program: Exospheric Mg II and Fe II in the Near-ultraviolet Transmission Spectrum of WASP-121b Using Jitter Decorrelation. The Astronomical Journal 158:91
- Sing et al. (2019b) Sing DK, Lavvas P, Ballester GE et al. (2019b) The Hubble Space Telescope PanCET Program: Exospheric Mg ii and Fe ii in the Near-ultraviolet Transmission Spectrum of WASP-121b Using Jitter Decorrelation. The Astronomical Journal 158(2):91
- Socrates (2013) Socrates A (2013) Relationship Between Thermal Tides and Radius Excess. arXiv:13044121 [astro-ph]
- Spiegel and Burrows (2013) Spiegel DS Burrows A (2013) Thermal Processes Governing Hot-Jupiter Radii. The Astrophysical Journal 772:76
- Steinrueck et al. (2021) Steinrueck ME, Showman AP, Lavvas P et al. (2021) 3D simulations of photochemical hazes in the atmosphere of hot Jupiter HD 189733b. Monthly Notices of the Royal Astronomical Society 504(2):2783–2799
- Sullivan et al. (2015) Sullivan PW, Winn JN, Berta-Thompson ZK et al. (2015) The Transiting Exoplanet Survey Satellite: Simulations of Planet Detections and Astrophysical False Positives. The Astrophysical Journal 809:77
- Thompson (1990) Thompson SL (1990) ANEOS analytic equations of state for shock physics codes input manual. Tech. Rep. SAND-89-2951, 6939284, Sandia National Laboratory, DOI 10.2172/6939284
- Thorngren and Fortney (2019) Thorngren D Fortney JJ (2019) Connecting Giant Planet Atmosphere and Interior Modeling: Constraints on Atmospheric Metal Enrichment. The Astrophysical Journal 874(2):L31
- Thorngren et al. (2019a) Thorngren D, Gao P Fortney JJ (2019a) The Intrinsic Temperature and Radiative-Convective Boundary Depth in the Atmospheres of Hot Jupiters. The Astrophysical Journal 884:L6
- Thorngren and Fortney (2018) Thorngren DP Fortney JJ (2018) Bayesian Analysis of Hot-Jupiter Radius Anomalies: Evidence for Ohmic Dissipation? The Astronomical Journal 155:214
- Thorngren et al. (2016) Thorngren DP, Fortney JJ, Murray-Clay RA Lopez ED (2016) The Mass-Metallicity Relation for Giant Planets. The Astrophysical Journal 831:64
- Thorngren et al. (2019b) Thorngren DP, Marley MS Fortney JJ (2019b) An Empirical Mass-Radius Relation for Cool Giant Planets. arXiv e-prints 1909:arXiv:1909.09,207
- Thorngren et al. (2021) Thorngren DP, Fortney JJ, Lopez ED, Berger TA Huber D (2021) Slow Cooling and Fast Reinflation for Hot Jupiters. The Astrophysical Journal 909:L16
- Thorngren et al. (2023) Thorngren DP, Lee EJ Lopez ED (2023) Removal of Hot Saturns in Mass-Radius Plane by Runaway Mass Loss. The Astrophysical Journal 945:L36
- Tremblin et al. (2017) Tremblin P, Chabrier G, Mayne NJ et al. (2017) Advection of Potential Temperature in the Atmosphere of Irradiated Exoplanets: A Robust Mechanism to Explain Radius Inflation. The Astrophysical Journal 841:30
- Turner et al. (2024) Turner JD, Grießmeier JM, Zarka P, Zhang X Mauduit E (2024) Follow-up LOFAR observations of the $ Boötis exoplanetary system. DOI 10.48550/ARXIV.2403.16392
- Valsecchi et al. (2014) Valsecchi F, Rasio FA Steffen JH (2014) From Hot Jupiters to Super-Earths via Roche Lobe Overflow. The Astrophysical Journal 793(1):L3
- Valsecchi et al. (2015) Valsecchi F, Rappaport S, Rasio FA, Marchant P Rogers LA (2015) Tidally-driven Roche-lobe Overflow of Hot Jupiters with MESA. The Astrophysical Journal 813:101
- Vazan et al. (2015) Vazan A, Helled R, Kovetz A Podolak M (2015) Convection and Mixing in Giant Planet Evolution. The Astrophysical Journal 803:32
- Vazan et al. (2016) Vazan A, Helled R, Podolak M Kovetz A (2016) The Evolution and Internal Structure of Jupiter and Saturn with Compositional Gradients. The Astrophysical Journal 829:118
- Vazan et al. (2018) Vazan A, Ormel CW Dominik C (2018) Effect of core cooling on the radius of sub-Neptune planets. Astronomy and Astrophysics 610:L1
- Wahl et al. (2017) Wahl SM, Hubbard WB, Militzer B et al. (2017) Comparing Jupiter interior structure models to Juno gravity measurements and the role of a dilute core. Geophysical Research Letters 44:4649–4659
- Wahl et al. (2021) Wahl SM, Thorngren D, Lu T Militzer B (2021) Tidal Response and Shape of Hot Jupiters. The Astrophysical Journal 921:105
- Wang et al. (2020) Wang HS, Morel T, Quanz SP Mojzsis SJ (2020) Europium as a lodestar: Diagnosis of radiogenic heat production in terrestrial exoplanets. arXiv:201004632 [astro-ph]
- Weber et al. (2018) Weber C, Erkaev NV, Ivanov VA et al. (2018) Supermassive Hot Jupiters Provide More Favourable Conditions for the Generation of Radio Emission via the Cyclotron Maser Instability - A Case Study Based on Tau Bootis b. arXiv:180711920 [astro-ph]
- Weiss et al. (2013) Weiss LM, Marcy GW, Rowe JF et al. (2013) The Mass of KOI-94d and a Relation for Planet Radius, Mass, and Incident Flux. The Astrophysical Journal 768:14
- Wong et al. (2021) Wong I, Shporer A, Zhou G et al. (2021) TOI-2109: An Ultrahot Gas Giant on a 16 hr Orbit. The Astronomical Journal 162(6):256
- Wu and Lithwick (2013) Wu Y Lithwick Y (2013) Ohmic Heating Suspends, Not Reverses, the Cooling Contraction of Hot Jupiters. The Astrophysical Journal 763:13
- Yadav and Thorngren (2017) Yadav RK Thorngren DP (2017) Estimating the Magnetic Field Strength in Hot Jupiters. The Astrophysical Journal Letters 849:L12
- Yee et al. (2021) Yee SW, Winn JN Hartman JD (2021) How Complete Are Surveys for Nearby Transiting Hot Jupiters? The Astronomical Journal 162(6):240
- Youdin and Mitchell (2010) Youdin AN Mitchell JL (2010) The Mechanical Greenhouse: Burial of Heat by Turbulence in Hot Jupiter Atmospheres. The Astrophysical Journal 721:1113–1126
- Zarka (1998) Zarka P (1998) Auroral radio emissions at the outer planets: Observations and theories. Journal of Geophysical Research: Planets 103(E9):20,159–20,194
- Zarka (2007) Zarka P (2007) Plasma interactions of exoplanets with their parent star and associated radio emissions. Planetary and Space Science 55(5):598–617
- Zhou et al. (2021) Zhou G, Quinn SN, Irwin J et al. (2021) Two Young Planetary Systems around Field Stars with Ages between 20 and 320 Myr from TESS. The Astronomical Journal 161:2